Italian Queue 2014 Next

Jun 15-16

Observer: E. Sani, A. Marchetti, M. Fumana
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: MODS/LUCI

Summary

Before the sunset heavy clouds over the dome.
At the sun set clouds at the horizon.
The sky is clear after the LUCI switch with low wind and seeing 0.7-1.0.
The LUCI detector has noise pattern related to problem with the readout (the reason is stil unknown). The pattern is even variable: sometimes there are just a cuople of stripes in top right quadrants, and some other all the quadrants are strongly affected.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard hz44 red grating -  
Standard hz44 dual grating -  
Bellazzini DDT HVC274.68+74 red grating 1h mostly within seeing specs
Beltran_01 A6 H2/Ks 50m/7.5m within seeing specs
Beltran_01 A1 H2/Ks 50m/7.5m within seeing specs
Telluric FS-148 H2/Ks    

Details

01:46 Bellazzini LS calibrations
  • (CalibScripts) grflats_red_LS60x5x1.2.cal 1r.0001-0005
  • (MODSeng) grpixflats_r.cal 1r.0006-0010
  • (MODSeng) grlamps_r.cal 1r.00011-0013

02:30 The TO opens the dome

BELLAZZINI DDT

03:13 Standard red grating hz44 1r.0016-0018
  • acqMODS hz44.acq
  • modsAlign -r mods1r.20140616.0015.fits
  • execMODS hz44.obs

03:35 Standard dual grating hz44 1r.0019-0021 1b.0001-0003
  • Cirrus
  • execMODS hz44_dual.obs

Wait for a darker sky

04:00 Slew to Bellazzini_46DDT
  • modsAlign -l mods1r.20140616.0022.fits mods1r.20140616.0023.fits target very faint
  • Try with a experimental version of modsAlign which subtract bias stripes but it doesn't work on LS masks
  • Use again the standard modsAlign Finding chart too small and too deep, more time is required to check alignment

04:21 Bellazzini_46DDT SCI scripts
  • execMODS HVC274.68+74_LS.acq 1r.0022-0024
  • Clouds saw on guide star monitor

04:30 execMODS HVC274.68+74_1.obs 1r.0025-0026
  • Due to thin cloud passing, we missed the guidestar for few minutes, the telescope was correctly tracking.
  • The average seeing on the GCS is 1.2 (sometimes up to 1.4)

05.27 We ask the TO to SWITCH TO LUCI

5:44 We are authorized to use LUCI. Already on Beltran target A6
  • executeLUCIscript A6_H2..img luci.20140616.0001
  • manual image to check the exp time on a single frame luci.20140616.0002
  • The script is not optimized, to much overhead to long exposures without switching on the sky.

06:18 executeLUCIscript A6_H2_DJT.img luci.20140616.0003-0039
  • the average seeing on the GCS (R band) is 0.8

08.05 executeLUCIscript A6_Ks_DJT.img luci.20140616.0040-0048
  • the average seeing on the GCS is 0.8

08.28 executeLUCIscript A1_H2_DJT.img luci.20140616.0049-0050

08.38 Slew to Cresci_8
  • the guider fails becausethe targets is too close to the Moon. We slew back to Beltran_01.

08.46 executeLUCIscript A1_H2_DJT.img luci.20140616.0051-0081
  • the average seeing on the GCS in between 0.7 and 0.9

10:15 * executeLUCIscript A1_Ks_DJT.img luci.20140616.0082-0092

10:48 executeLUCIscript FS-148_H2.img luci.20140616.0093-0097

11:00 executeLUCIscript FS-148_Ks.img luci.20140616.0098-0102

11:10 Switch to calibrations
  • Sky Ks flats off (low counts) luci.20140616.0103-0202
  • Sky Ks flats off (high counts) luci.20140616.0215-0024
  • Sky H2 flats on (high counts) luci.20140616.0227-0236

12:11 The TO closes the enclosure
  • we run dark frames for the Beltran and Cresci programs.

-- %USERSIG{MarcoFumana - 2014-06-16}%
Topic revision: r6 - 17 Jun 2014, MarcoFumana
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback